Three cards are selected at random without replacement from a well-shuffled deck of 52 playing cards. Find the probability that three cards of the same suit are drawn. (Round your answer to four decimal places.)
Found 2 solutions by checkley77, NightMareZ:
Answer by checkley77(12844)

13/52*12/51*11/50=1,716/132,600=.0129 ANS.OR 1.29%.

Answer by NightMareZ(1)

There are 4 suit ,, hearts , Diamonds , clubs , Spades
then the 3 cards may be from the 1st OR 2nd OR 3rd OR 4th
p(A)=4*(13C3)/(52C3)=0.0517 Or 5.17%
@
Mohamed ,, :)